OpenStudy (dido625):

Set up a definite integral that represents the volume obtained by rotating the region between the curve y^2=x and y^2=2(x-1) about the lines below. (a) Rotate about the y-axis (b) Rotate about the line y=3

OpenStudy (dido625):

Normally, I am okay with elementary calculus but I always struggle with this part. Any guidance?

OpenStudy (nincompoop):

first, try and draw

OpenStudy (dido625):

I already drew out the region. Looks like a washer to me.

OpenStudy (nincompoop):

how do you determine between washer and shell method?

OpenStudy (dido625):

Inner and outer radius for Washer and cylinders.

OpenStudy (nincompoop):

or even disk

OpenStudy (dido625):

Yea. Disk too.

OpenStudy (nincompoop):

so if you already know that it is washer, then you just have to use the formula

OpenStudy (dido625):

I know. I'm just having writing it out xD xD .Hold on.

OpenStudy (dido625):

\[\int\limits_{a}^{b}\pi(x^2-(2(x-1))^2 dx\]

OpenStudy (dido625):

Is that it? I can easily find a and by setting the two equations equal to each other.

OpenStudy (nincompoop):

then do it

OpenStudy (dido625):

\[-\sqrt{2},\sqrt{2}\] ?